### v1

# SAINT: Attention-Based Modeling of Sub-Action Dependencies in Multi-Action Policies

Abstract
The combinatorial structure of many real-world action spaces leads to exponential growth in the number of possible actions, limiting the effectiveness of conventional reinforcement learning algorithms. Recent approaches for combinatorial action spaces impose factorized or sequential structures over sub-actions, failing to capture complex joint behavior. We introduce the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that represents multi-component actions as unordered sets and models their dependencies via self-attention conditioned on the global state. SAINT is permutation-invariant, sample-efficient, and compatible with standard policy optimization algorithms. In 15 distinct combinatorial environments across three task domains, including environments with nearly 17 million joint actions, SAINT consistently outperforms strong baselines. [^1] Code is available at https://github.com/matthewlanders/SAINT
1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a range of domains, primarily through methods designed for either small, discrete action spaces [15, 28, 41, 27], such as those in classic games, or continuous control [11, 12, 22, 33] typical of robotics. However, many real-world problems fall between these extremes, involving large, discrete combinatorial action spaces defined as Cartesian products of multiple subspaces, with each joint action comprising several coordinated sub-actions: $\mathbf{a}=(a_{1},\dots,a_{A})$. Such settings arise in critical applications including traffic signal control [32], drug selection in healthcare [35], and job scheduling in multi-resource compute clusters [24].
Traditional RL methods typically model discrete action spaces using a flat categorical policy. In combinatorial settings, this becomes intractable as the number of possible actions grows exponentially with the number of sub-actions, scaling as $\prod_{d=1}^{A}m_{d}$ for $A$ sub-action dimensions, where $m_{d}$ denotes the number of discrete choices in the $d$-th dimension. To mitigate this combinatorial explosion, existing approaches [8, 31, 38, 47] often rely on simplifying assumptions that constrain the representational capacity of the policy class. A common strategy assumes conditional independence among sub-actions given the state, factorizing the policy as $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s)$ [31, 38]. While this provides computational efficiency, it fails to capture cooperative or antagonistic interactions between sub-actions, limiting the policy’s ability to express complex joint behaviors. An alternative is to impose a fixed autoregressive order, modeling the policy as $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s,a_{<i})$ [47]. Although more expressive, this approach introduces an arbitrary sequence over sub-actions — even in domains where such orderings are not meaningful — breaking permutation invariance and potentially impairing learning when the imposed order misaligns with the true dependency structure.
Many real-world tasks violate both the independence and fixed-order assumptions. For example, in healthcare, drug combinations often exhibit complex interaction effects: two drugs may be beneficial individually but harmful together. Accurately modeling such dependencies requires policy architectures that can capture joint interactions without relying on conditional independence or arbitrary sequencing assumptions. This motivates the development of more flexible, permutation-invariant models for combinatorial action spaces.
We propose the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that models multi-component actions as unordered sets of sub-actions. SAINT combines principles from permutation-invariant set processing with the Transformer’s self-attention mechanism [21, 43] to capture high-order dependencies among sub-actions. To incorporate global state information, SAINT first applies Feature-wise Linear Modulation (FiLM) [30] to the initial sub-action embeddings. These state-conditioned embeddings are then processed by Transformer blocks to model context-dependent interactions in parallel. Finally, each sub-action’s representation is decoded separately, preserving tractability without assuming independence and maintaining permutation invariance.
We empirically evaluate SAINT on challenging benchmark tasks that exhibit both primarily state-independent and primarily state-dependent sub-action dependencies. Our results show that SAINT significantly outperforms strong baselines, including those with independent and autoregressive factorizations, across action spaces ranging from hundreds to over 16 million discrete actions. Targeted ablations validate the effectiveness of SAINT’s pre-attention FiLM-based state conditioning and quantify the computational overhead introduced by its attention mechanism. Together, these findings establish SAINT’s ability to efficiently learn coordinated behaviors in complex combinatorial action spaces.
2 Related Work
Combinatorial Action Spaces
Combinatorial action spaces arise naturally in many sequential decision-making problems, including text-based interaction, games, and resource allocation. Prior approaches have introduced task-specific architectures or assumptions tailored to these settings [4, 5, 7, 13, 14, 29, 44]. Others exploit problem-specific structure, such as the amenability to gradually expanding curricula, to improve sample efficiency [10]. While effective in context, such methods typically lack generality and require significant manual design effort. In contrast, SAINT provides a domain-agnostic architecture that models sub-action dependencies through permutation-invariant self-attention, offering greater flexibility across structured decision problems. A parallel body of work addresses continuous control problems by discretizing the action space [2, 25, 36, 40], which contrasts with our focus on inherently discrete action spaces with combinatorial structure. Recent work in offline RL has explored learning in combinatorial action spaces under dataset constraints [3, 20, 35]; however, these approaches address challenges that differ from those in online learning.
A number of general-purpose architectures have been developed to scale RL to high-dimensional, combinatorial action spaces. One common strategy is to factorize the action space, treating each sub-action as conditionally independent given the state. This reduces the action space size from exponential to linear in the number of sub-actions [31, 38]. However, the independence assumption limits the expressiveness of the policy and breaks down in settings where sub-actions strongly interact. Autoregressive (AR) models [47] instead decompose the joint action distribution into a sequence of conditional distributions, allowing the policy to model dependencies among sub-actions. Yet, this introduces an arbitrary ordering over sub-actions — even when no natural order exists — violating permutation invariance and potentially impairing learning when the imposed structure misaligns with the task. The Wolpertinger architecture (Wol-DDPG) [8] embeds discrete actions into a continuous space and uses approximate nearest-neighbor search to retrieve a subset of candidate actions before value-based refinement. While scalable, this method performs poorly when the structure of the action space is inherently unordered [5], a limitation we verify experimentally.
Transformers in Reinforcement Learning
The success of the Transformer architecture [43] in sequence modeling has led to its adoption in RL, most notably through methods like Decision Transformer (DT) [6] and Trajectory Transformer (TT) [17]. These approaches use Transformer decoders to autoregressively predict actions based on past trajectories of states, actions, and returns, showing strong performance in offline RL.
Transformers have also been used in RL settings beyond trajectory modeling [1]. DTQN [9] applies a Transformer decoder for Q-learning in partially observable environments, while other work uses them for predictive world modeling [26, 34] and high-dimensional state representation learning [18, 19]. Attention mechanisms over state representations have also been shown to improve reasoning in complex environments. For example, multi-head attention in centralized critics enables agents to selectively aggregate information from others, improving stability and sample efficiency [16], while self-attention over learned entity embeddings supports relational inference and generalization [45].
In contrast to prior approaches that apply attention to model relations among external entities, agents, or trajectories, SAINT leverages self-attention over sub-actions within a single timestep. Inspired by the Transformer’s success in set processing [21], SAINT treats sub-actions as an unordered set and applies permutation-equivariant self-attention to model their interactions, enabling the policy to capture complex intra-action dependencies critical for combinatorial control.
3 Preliminaries
Combinatorial Action Spaces
We consider RL problems formalized as a finite-horizon Markov Decision Process, defined by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,r,\gamma,\mu\rangle$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, and $\mu$ the initial state distribution. A policy $\pi$ maps states to distributions over actions, $\pi:\mathcal{S}\rightarrow\mathbb{P}(\mathcal{A})$, thereby governing the agent’s behavior over episodes of length $H$.
In this work we assume actions have an explicit compositional structure. Specifically, the action space is a product of sub-action domains, $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{A}$, where each sub-action space $\mathcal{A}_{i}$ can be either discrete or continuous. A single action thus comprises $A$ sub-decisions, $\mathbf{a}=(a_{1},\dots,a_{A})$, with each $a_{i}\in\mathcal{A}_{i}$. This representation gives rise to high-dimensional action spaces with potentially complex dependencies among sub-actions.
Attention
Attention is a general computational primitive that allows a model to selectively aggregate information from a set of inputs based on learned relevance scores [43]. Formally, given a set of queries $Q\in\mathbb{R}^{n_{q}\times d}$, keys $K\in\mathbb{R}^{n_{k}\times d}$, and values $V\in\mathbb{R}^{n_{k}\times d}$, the scaled dot-product attention computes output $\mathrm{Attn}(Q,K,V)=\text{softmax}(QK^{\top}/\sqrt{d})V$. Multi-head self-attention extends this by learning multiple independent projections and aggregating their outputs, enabling the model to capture different types of interactions in parallel. Transformers, which stack layers of multi-head self-attention with feedforward components, have been widely adopted in domains requiring flexible modeling of structured dependencies.
4 Sub-Action Interaction Network using Transformers (SAINT)
Figure 1: Overview of SAINT. Initial sub-action embeddings are conditioned on the state $s$ using FiLM. Stacked Transformer blocks then model dependencies among sub-actions. The resulting context-aware representations are passed to independent Decision MLPs, which output per-sub-action policy distributions used for factorized sampling.
Prior RL methods for learning in combinatorial action spaces typically represent sub-actions either as conditionally independent or in a fixed autoregressive sequence. These approaches limit policy expressiveness and are ill-suited for domains where sub-actions naturally form an unordered set, such as in traffic signal control. To address these limitations, we propose the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that models multi-component actions as unordered sets of $A$ sub-actions. SAINT first applies Feature-wise Linear Modulation (FiLM) [30] to inject global state information into sub-action embeddings drawn from a learned embedding table. These state-aware embeddings are then processed by multi-head Transformer blocks, enabling the model to capture higher-order dependencies among sub-actions. The resulting representations are passed through individual MLP decision networks to parameterize per-sub-action distributions, from which actions are sampled independently. This enables efficient action construction while preserving the ability to model complex interactions. To our knowledge, SAINT is the first policy architecture to model complex sub-action dependencies in combinatorial action spaces while preserving permutation equivariance. An overview of the SAINT architecture is shown in Figure 1.
Given an environment state $\mathbf{s}\in\mathbb{R}^{d_{s}}$, SAINT begins by assigning each sub-action $i\in\{1,\dots,A\}$ a learnable embedding from a table $\mathrm{Embed}\in\mathbb{R}^{A\times d}$, yielding a base representation $\mathbf{e}_{i}=\mathrm{Embed}(i)\in\mathbb{R}^{d}$. Each sub-action index $i$ is thus mapped to a learnable $d$-dimensional vector, regardless of the original cardinality of $\mathcal{A}_{i}$, providing a shared embedding space that enables uniform processing by the subsequent Transformer layers.
To condition the sub-action representations on the global state, we apply FiLM to the initial embeddings $\mathbf{e}_{i}$. A shared MLP $g:\mathbb{R}^{d_{s}}\to\mathbb{R}^{2d}$ maps the state $\mathbf{s}$ to FiLM parameters $(\boldsymbol{\gamma},\boldsymbol{\beta})=g(\mathbf{s})$, which are then used to apply a sub-action-wise affine transformation:
$\tilde{\mathbf{e}}_{i}=\boldsymbol{\gamma}\odot\mathbf{e}_{i}+\boldsymbol{\beta}\;.$
Following FiLM-based conditioning, the matrix of state-aware sub-action representations, $\tilde{\mathbf{E}}\in\mathbb{R}^{A\times d}$, is processed by a stack of $L$ identical Transformer blocks, omitting positional encodings to preserve permutation equivariance. Letting $\mathbf{X}^{(0)}=\tilde{\mathbf{E}}$, each block $\ell=1,\dots,L$ first computes multi-head self-attention over the sub-actions. Specifically, it constructs queries, keys, and values via learned projections: $\mathbf{Q},\mathbf{K},\mathbf{V}=\mathbf{X}^{(\ell-1)}W^{Q},\mathbf{X}^{(\ell-1)}W^{K},\mathbf{X}^{(\ell-1)}W^{V}$, and applies scaled dot-product attention to model interactions among sub-actions. The output of the attention layer is passed through a position-wise feed-forward network (FFN) applied independently to each sub-action embedding. Each sub-layer, attention and FFN, is followed by residual connections and layer normalization, as in standard Transformer architectures. This mechanism enables the policy to model state-conditioned dependencies between sub-actions while maintaining permutation equivariance.
The final output $\mathbf{X}=\mathbf{X}^{(L)}\in\mathbb{R}^{A\times d}$ provides a representation $\mathbf{x}_{i}$ for each sub-action $i$ that integrates global state information from FiLM conditioning and inter-sub-action dependencies modeled via self-attention. A sub-action-specific decision MLP $f_{i}:\mathbb{R}^{d}\to\mathbb{R}^{P_{i}}$ then maps $\mathbf{x}_{i}$ to the parameters of a distribution $\pi_{i}$ over the sub-action space $\mathcal{A}_{i}$. For example, if $\mathcal{A}_{i}$ is discrete with $K_{i}$ possible actions, $f_{i}$ outputs $P_{i}=K_{i}$ logits for a categorical distribution $\pi_{i}(a_{i}\mid\mathbf{s})=\mathrm{Categorical}(f_{i}(\mathbf{x}_{i}))$. If $\mathcal{A}_{i}$ is continuous, $f_{i}$ may output $P_{i}=2$ values — the mean and variance of a Gaussian distribution.
Because each sub-action representation encodes state and interaction context, the policy can be factorized across sub-actions without loss of modeling capacity:
$\log\pi(\mathbf{a}\mid\mathbf{s})=\sum_{i=1}^{A}\log\pi_{i}(a_{i}\mid\mathbf{s})\;.$
This factorization is necessary in combinatorial action spaces, where evaluating or optimizing over the full joint distribution $P(a_{1},\dots,a_{A}\mid\mathbf{s})$ is typically infeasible.
A separate value network $V:\mathbb{R}^{d_{s}}\to\mathbb{R}$ provides a scalar estimate of the state value. In our implementation, $V$ is an MLP that conditions only on $\mathbf{s}$ and does not share parameters with the policy network, though shared encoders could be used for efficiency if desired. SAINT is agnostic to the choice of learning algorithm and can be integrated with standard policy gradient and actor-critic methods, such as PPO [33] and A2C [27], using their corresponding training objectives.
By combining FiLM-based state conditioning with permutation-equivariant self-attention, SAINT provides a unified architecture for learning policies in combinatorial action spaces. Its design enables rich, state-dependent interaction modeling among sub-actions while maintaining scalability through independent decoding. The resulting factorized policy is both expressive and tractable, capable of operating in high-dimensional action spaces without assuming independence or imposing arbitrary ordering. Together with a modular value function and compatibility with standard RL algorithms, SAINT offers a flexible and general-purpose framework for structured decision-making. Pseudocode for action selection using SAINT is provided in Algorithm 1.
Algorithm 1 Sub-Action Interaction Network using Transformers (SAINT)
1:Construct initial embedding matrix $\mathbf{E}\leftarrow[\text{Embed}(1),\dots,\text{Embed}(A)]^{\top}$triangleright $\mathbf{E}\in\mathbb{R}^{A\times d}$
2:Compute FiLM parameters from state $(\boldsymbol{\gamma},\boldsymbol{\beta})\leftarrow g(\mathbf{s})$
3:Apply FiLM conditioning element-wise $\tilde{\mathbf{E}}\leftarrow\boldsymbol{\gamma}\odot\mathbf{E}+\boldsymbol{\beta}$
4:Process through Transformer blocks $\mathbf{X}\leftarrow\text{TransformerBlocks}(\tilde{\mathbf{E}})$triangleright $\mathbf{X}\in\mathbb{R}^{A\times d}$
5:Initialize action vector $\mathbf{a}=(\dots)$
6:Initialize total log-probability $\ell\leftarrow 0$
7:for $i=1$ to $A$ do $\triangleright$ Independent decoding per sub-action
8:     Extract context-aware embedding for sub-action $\mathbf{x}_{i}\leftarrow\mathbf{X}[i,:]^{\top}$
9:     Compute distribution parameters using decision network $\text{params}_{i}\leftarrow f_{i}(\mathbf{x}_{i})$
10:     Define distribution $\pi_{i}\leftarrow\text{Distribution}(\mathcal{A}_{i},\text{params}_{i})$triangleright e.g., Categorical
11:     Sample action $a_{i}\sim\pi_{i}$
12:     Accumulate log-probability $\ell\leftarrow\ell+\text{log-prob}(a_{i};\pi_{i})$
13:     Store sampled sub-action $\mathbf{a}[i]\leftarrow a_{i}$
14:end for
15:return Return joint action $\mathbf{a}=(a_{1}..a_{A})$ and total log-prob $\ell$
5 Experimental Evaluation
In our experiments, we address three primary research questions: (1) How effective is SAINT in environments with primarily state-independent sub-action dependencies? (2) How effective is SAINT in environments with primarily state-dependent sub-action dependencies? (3) How effective is SAINT in environments with weak sub-action dependencies but complex dynamics? We address these questions in Sections 5.15.2, and 5.3, respectively. In Section 5.4, we present ablation studies that compare FiLM-based conditioning to alternative state injection strategies and measure the computational cost of SAINT’s architecture.
We compare SAINT to four baselines: (1) a pure factorization approach with independent sub-action policies, (2) an autoregressive model [47], (3) Wol-DDPG [8], and (4) a standard RL algorithm selected based on best empirical performance in each environment (PPO in Sections 5.1 and 5.3; A2C in Section 5.2). All reported results are averaged over five random seeds.
5.1 State-Independent Sub-Action Dependencies
To evaluate SAINT in environments where sub-action dependencies are primarily state-independent, we use the CityFlow traffic control benchmark [46], where each action corresponds to simultaneous phase decisions across multiple intersections. While coordination is necessary to achieve global traffic efficiency, the structure of these dependencies remains largely unchanged across states.
We consider two configurations: (1) CityFlow-Linear, where three traffic signals are arranged in a row, yielding 729 possible joint signal combinations; and (2) CityFlow-Irregular, with four closely spaced, asymmetric intersections and varying road capacities, resulting in 375 valid joint action combinations. These networks are illustrated in Appendix A.1.
We adopt PPO as the standard RL algorithm in this setting. SAINT, the pure factorized baseline, and the AR baseline, use the same PPO implementation and hyperparameters, isolating the impact of architectural differences. Wol-DDPG performs poorly in this environment and is excluded from the main learning curves; full training curves, including Wol-DDPG, are included in Appendix A.2. Wol-DDPG’s ineffectiveness aligns with prior observations that it is ill-suited to environments with unordered sub-actions [5].
(a) CityFlow-Linear
(b) CityFlow-Irregular
Figure 2: Learning curves show that SAINT outperforms all baselines in both learning speed and final reward, across both the CityFlow-Linear and CityFlow-Irregular environments. In CityFlow-Irregular, SAINT’s relative advantage widens as coordination demands increase.
As can be seen in Figure 2, SAINT learns faster and achieves higher final performance than all baselines in both environments. The gains are especially pronounced in CityFlow-Irregular, where the increased coordination complexity makes the limitations of factorized and sequential approaches more apparent. Notably, PPO, which treats the entire joint action space monolithically, fails to learn a policy approaching the performance of SAINT in either setting.
5.2 State-Dependent Sub-Action Dependencies
Next, we evaluate SAINT in environments where sub-action dependencies are strongly state-dependent. In these domains, the relationships among sub-actions vary significantly with the environment state, requiring policies to model context-sensitive joint decisions. To study this setting, we use the Combinatorial Navigation Environment (CoNE) [20], a configurable high-dimensional control domain designed to evaluate policy architectures under large, discrete, and structured action spaces.
The agent begins at a fixed origin $s_{0}$ and must reach a predefined goal $g$. At each timestep, it selects a joint action by activating multiple discrete sub-actions, each corresponding to movement along a distinct dimension of the environment. These sub-actions are executed in parallel to produce a single composite transition. The agent receives a reward $r=-\rho(s,g)$ based on the Euclidean distance to the goal. Episodes terminate either upon reaching the goal (reward $+10$) or entering a terminal failure state (pit), which incurs a penalty of $r=-10\cdot\rho(s_{0},g)$ to discourage reward hacking through early failure.
In CoNE, both the action and state spaces grow exponentially with dimensionality: the number of joint actions scales as $|\mathcal{A}|=2^{2D}$, and the number of states as $|\mathcal{S}|=M^{D}$ in a $D$-dimensional environment with $M$ positions per axis. In our largest setting, the environment comprises over 200 million states and nearly 17 million joint actions per state. Beyond scale, CoNE introduces strong sub-action dependencies: some combinations enable efficient movement, others cancel out, and some lead to catastrophic failure. Crucially, these sub-action interactions are highly state-dependent; a combination that is optimal in one state may lead to a pit in another, making effective decision-making highly sensitive to global context. CoNE is highly configurable, allowing us to systematically vary the number of dimensions and pit density to assess the impact of increasing action space size and sub-action dependence.
We adopt A2C as the standard RL algorithm in this setting. SAINT, the pure factorized baseline, and the AR baseline use identical A2C implementations to ensure a controlled comparison.
Varying Dimensionality
To evaluate SAINT’s effectiveness as the number of possible actions increases, we scale the dimensionality of CoNE from 7 (yielding over 16 thousand possible action combinations) to 12 (with nearly 17 million combinations). In CoNE environments without pits, the agent can learn a trivial policy, as the optimal solution involves selecting the same action in every state. Thus, to introduce meaningful complexity and prevent this degenerate behavior, we place pits in 25% of interior states.
$|\mathcal{A}|$ SAINT Factored AR Wol-DDPG A2C
[…]
Table 1: Performance in CoNE as action dimensionality increases. SAINT consistently achieves the highest reward across all action space sizes. Factorized and autoregressive baselines plateau at substantially lower reward levels, while Wol-DDPG and A2C fail to learn viable policies.
The results in Table 1 show that SAINT maintains strong performance as action dimensionality increases, significantly outperforming all baselines at every scale. While the factorized and autoregressive baselines achieve modest performance in lower dimensions, their performance degrades or plateaus as the number of sub-actions grows. SAINT’s relative advantage increases in the largest settings, where it maintains low variance and stable performance. Wol-DDPG and A2C perform poorly throughout, likely due to their inability to scale to large, unordered action spaces. Note that A2C is omitted at the highest dimensionality due to the computational intractability of modeling the full joint action space with a flat categorical policy. Full learning curves are provided in Appendix B.2.
Varying Dependence
To assess SAINT’s robustness to different levels of sub-action dependence, we incrementally increased the number of pits in the 12-dimensional CoNE environment. Higher pit densities impose stronger coordination requirements, as more sub-action combinations must be carefully selected to avoid pits. To ensure that a valid path from the start state to the goal always exists, pits were placed only in interior (non-boundary) states. We generate environments with 10%, 25%, 50%, 75%, and 100% of interior states occupied by pits. Note that even in the 100% setting, all boundary states remain pit-free, guaranteeing the existence of at least one (possibly inefficient) path to the goal region. We exclude A2C from this experiment due to the computational intractability of modeling such a large discrete action space (nearly 17 million actions) with a flat categorical distribution.
The results in Table 2 demonstrate that SAINT is more robust to increasing sub-action dependence than all baselines. As pit density increases from 10% to 100%, SAINT maintains high performance with low variance, while the factorized and autoregressive baselines generally degrade. Wol-DDPG was unable to learn meaningful policies at any pit density. These results highlight SAINT’s ability to capture complex, context-sensitive dependencies between sub-actions that are critical in many real-world combinatorial environments. Full learning curves for these results are provided in Appendix B.3.
Pit % SAINT Factored AR Wol-DDPG
[…]
Table 2: Performance in 12-D CoNE ($\sim 17$M actions) as sub-action dependence increases (controlled via pit density). SAINT consistently achieves the highest rewards across all settings and remains robust even as the sub-actions become highly dependent. Other methods degrade more rapidly, especially the factorized baseline. Wol-DDPG failed to learn meaningful policies in this setting.
5.3 Weak Sub-Action Dependencies with Complex Dynamics
To evaluate SAINT in environments where sub-action dependencies are relatively weak but the underlying dynamics are complex, we consider discretized variants of the HalfCheetah, Hopper, and Walker2D MuJoCo locomotion tasks [39]. In these environments, each continuous joint control signal is discretized into 11 bins, yielding large, structured action spaces while retaining the rich temporal and physical dynamics of the original tasks. Although the discretized action spaces are combinatorially large, prior work [3] suggests that the dependencies among sub-actions are relatively weak in these domains. This setting thus serves as a testbed for assessing whether SAINT’s architectural benefits persist even when strong coordination between sub-actions is not required.
Given the limitations of PPO and Wol-DDPG identified in Sections 5.1 and 5.2, we limit our comparison to the pure factorized and autoregressive baselines, using identical PPO implementations to ensure a controlled evaluation.
(a) HalfCheetah
(b) Hopper
(c) Walker2D
Figure 3: Performance in discretized MuJoCo environments. While results are similar across methods in HalfCheetah, SAINT outperforms factorized and autoregressive baselines in Hopper and Walker2D, demonstrating its ability to handle complex action spaces even when sub-action dependencies are relatively weak.
As shown in Figure 3, SAINT performs comparably to the factorized and autoregressive baselines in HalfCheetah. In Hopper and Walker2D, SAINT achieves faster learning and higher final returns, indicating that modeling sub-action structure can provide advantages even in settings with relatively weak dependencies.
5.4 Ablations
To assess the effect of SAINT’s architecture, we conduct targeted ablation studies. First, we compare SAINT’s pre-attention FiLM-based state conditioning to several alternative strategies for injecting state information. Second, we analyze the computational cost of SAINT relative to a pure factorization baseline, quantifying the trade-off between representational capacity and efficiency. These evaluations are performed in the CityFlow-Irregular environment and in the 10-dimensional CoNE setting with pits occupying 25% of interior states.
State Conditioning
We compare SAINT’s FiLM-based state conditioning to several alternative mechanisms for injecting global state information. As shown in Appendix C, FiLM yields higher final performance and more stable training in CityFlow, and performs at least as well as other methods in CoNE. These results suggest that although several conditioning mechanisms are effective, FiLM provides a performance advantage and may promote more stable training dynamics.
Attention Cost
To isolate the computational overhead of modeling sub-action dependencies via self-attention, we compare SAINT’s runtime to that of the pure factorization baseline. We also evaluate a variant of SAINT, called SAINT-IP, that replaces standard self-attention with an inducing point mechanism [21], which approximates full attention using a fixed set of learned summary vectors. This technique reduces the quadratic cost of attention by attending first from the inducing points to the inputs, and then from the inputs back to the summaries. All experiments were conducted on a single NVIDIA A40 GPU using Python 3.9 and PyTorch 2.6. We report wall-clock time per training episode in seconds, averaged over 5 runs.
SAINT SAINT-IP Factored
CityFlow
[…]
CoNE
[…]
Table 3: Total wall-clock training time and time to reach the factored baseline’s final performance, both measured in seconds. "Total training time" reflects the full duration of training, while "time to factored performance" measures how quickly each method reaches that baseline.
As shown in Table 3, SAINT incurs higher total training time than the pure factorization baseline, reflecting the additional cost of modeling sub-action dependencies via self-attention. SAINT-IP further increases this overhead due to the constant-factor cost of the Induced Set Attention Block (ISAB), which performs two full attention passes per layer. Standard self-attention (used in SAINT), by contrast, requires only one. Although the number of sub-actions in our environments is nontrivial, it remains small relative to domains like large language modeling or 3-D vision, for which attention-based inducing points were originally designed [21]. Consequently, the quadratic cost of full self-attention is not prohibitive, and ISAB’s asymptotic advantage does not yield runtime benefits in practice.
Although self-attention introduces per-step overhead, both SAINT and SAINT-IP reach the factored baseline’s asymptotic performance significantly faster. In CityFlow, they do so in less than one-third of the total training time; in CoNE, SAINT reaches this level nearly 30% faster. These results underscore SAINT’s practical efficiency: while modeling dependencies incurs additional cost, it leads to faster acquisition of strong policies. As shown in Appendix D, SAINT-IP exhibits some degradation in final performance relative to SAINT in CityFlow, but still outperforms the pure factorization baseline in both environments.
6 Discussion and Conclusion
We introduce SAINT, a Transformer-based policy architecture for combinatorial action spaces with interdependent sub-actions. Unlike existing approaches that assume independence or impose fixed autoregressive orderings, SAINT treats actions as unordered sets and models their structure using permutation-invariant self-attention. A FiLM-based global conditioning mechanism allows SAINT to incorporate state context efficiently, enabling it to capture rich, state-dependent sub-action interactions without sacrificing tractability.
While SAINT demonstrates strong performance and scalability in complex combinatorial action spaces, its limitations present opportunities for future work. First, the self-attention mechanism, while effective for modeling sub-action dependencies, introduces higher per-step computational cost than purely factorized baselines. Although Section 5.4 shows this is often offset by improved sample efficiency and faster convergence, lighter-weight attention variants (e.g., sparse attention or optimized ISAB for large $A$) could benefit resource-constrained settings or extremely high-dimensional action spaces. Second, our formulation assumes a fully observable MDP. While FiLM-based state conditioning enables adaptation to different contexts, handling significant partial observability may require incorporating memory modules or recurrent structures prior to the FiLM layer. Third, SAINT relies on the FiLM network to inject state information into sub-action embeddings. Although ablations show this to be effective, performance may be sensitive to the complexity of state-action relationships and the capacity of the conditioning network. Finally, while SAINT is designed as a general-purpose architecture and performs well across domains (e.g., traffic, navigation, locomotion), entirely novel tasks with different sub-action structures may require tuning. Future work could explore more robust generalization or meta-learning approaches for rapid adaptation to new combinatorial tasks.
Despite these limitations, extensive experiments show that SAINT consistently outperforms strong baselines in both learning speed and final performance, scaling to action spaces with nearly 17 million joint actions. Ablations confirm the effectiveness of our state injection strategy, and runtime analysis shows that the cost of modeling sub-action dependencies is offset by faster attainment of performant policies.
Appendix A Learning in CityFlow
A.1 Environmental Setup
(a) CityFlow-Linear
(b) CityFlow-Irregular
Figure 4: Visualizations of the two CityFlow traffic control configurations used in our experiments. CityFlow-Linear (Figure 4(a)) has three intersections arranged in a row, yielding 729 possible joint actions. CityFlow-Irregular (Figure 4(b)) has 375 joint actions but exhibits greater coordination demands and more diverse traffic interactions.
In both CityFlow-Linear (Figure 4(a)) and CityFlow-Irregular (Figure 4(b)), the state is represented as a flat integer vector, in which each value indicates the number of waiting vehicles on an incoming lane and its paired outgoing lane. The reward at each step is the negative of the average "pressure" across intersections, where an intersection’s pressure is defined as the absolute difference between its total incoming and outgoing vehicle counts. Pressure is a standard metric in traffic signal control literature, commonly used to quantify imbalance in intersection flow [23, 42].
A.2 Learning Curves Including Wol-DDPG
(a) CityFlow-Linear
(b) CityFlow-Irregular
Figure 5: Full learning curves for all baselines in CityFlow, including Wol-DDPG. Wol-DDPG performs poorly, consistent with its known limitations in environments with unordered sub-actions.
Figure 5 presents the full training curves for all methods in the CityFlow environments, including Wol-DDPG. As noted in Section 5.1, Wol-DDPG consistently underperforms relative to other methods. This poor performance is consistent with prior findings [5], which show that Wol-DDPG is ill-suited to settings with unordered sub-actions. Learning curves excluding Wol-DDPG are presented in Figure 2.
Appendix B CoNE Learning Curves
B.1 Environmental Setup
The Combinatorial Navigation Environment (CoNE) [20] is designed to evaluate RL algorithms in settings with high-dimensional, combinatorial action spaces and strong, state-dependent sub-action dependencies. In CoNE, actions are formed by simultaneously selecting discrete sub-actions, each specifying movement along a different dimension. These sub-actions are executed in parallel to produce a composite transition, which may advance the agent toward the goal or result in failure by entering a pit.
CoNE supports scaling along two axes, action dimensionality and pit density. As the number of dimensions increases, both the state and action spaces grow exponentially; our largest configuration contains over 200 million states and nearly 17 million discrete joint actions per state. In CoNE, sub-action interactions are complex: some combinations are efficient, others cancel each other out, and many must be avoided. These dependencies are highly state-sensitive, requiring effective decision-making to account for both structure and context.
To our knowledge, no other existing benchmarks offer the combination of large-scale action spaces and tunable sub-action dependencies found in CoNE. Popular environments such as the DeepMind Control Suite [37], for example, lack meaningful sub-action interactions [3].
B.2 Varying Dimensionality
Figure 6: Learning curves in CoNE environments as the number of sub-action dimensions increases from 7 to 12 (corresponding to joint action spaces ranging from $\sim 16$k to $\sim 17$M actions). SAINT consistently achieves higher final rewards than all baselines, with its advantage widening in higher-dimensional settings. Factorized and autoregressive baselines struggle to scale beyond moderate dimensions, while Wol-DDPG and A2C fail to learn meaningful policies across all tasks. Results are averaged over 5 seeds; shaded regions indicate one standard deviation.
Figure 6 provides the full learning curves corresponding to the results in Table 1, which reports performance in CoNE as the number of sub-action dimensions increases. As dimensionality grows, the joint action space expands exponentially — from roughly 16 thousand to nearly 17 million possible joint actions.
Across all settings, SAINT consistently outperforms baselines. Notably, SAINT maintains stable learning dynamics and low variance even at the largest scales, whereas factorized and autoregressive baselines generally plateau. Wol-DDPG and A2C fail to learn viable policies in any configuration, highlighting their inability to handle large, unordered combinatorial action spaces. These results underscore SAINT’s scalability and its robustness to increasing combinatorial complexity.
B.3 Varying Dependence
Figure 7: Learning curves in the 12-D CoNE environment as pit density increases from 10% to 100%, inducing progressively stronger sub-action dependencies. SAINT consistently outperforms all baselines, maintaining stable performance even as coordination requirements become increasingly stringent. Factored and autoregressive baselines generally plateau, while Wol-DDPG fails to learn meaningful policies. Results are averaged over 5 seeds; shaded regions denote one standard deviation.
Figure 7 shows the full learning curves corresponding to the results in Table 2, which reports performance in the 12-D CoNE environment as sub-action dependence increases via pit density. As more interior states are occupied by pits, successful navigation requires greater coordination among sub-actions to avoid failure states.
SAINT maintains stable learning and strong final performance across all pit densities, even as coordination requirements grow substantially. Factorized policies degrade, while autoregressive policies consistently underperform relative to SAINT. Wol-DDPG fails to make progress in any environment. A2C was excluded from this experiment due to the computational intractability of modeling such a large discrete action space (nearly 17 million actions) with a flat categorical distribution. These results highlight SAINT’s robustness to state-dependent sub-action dependencies.
Appendix C State Conditioning
We compare SAINT’s pre-attention FiLM-based state conditioning to four alternative mechanisms: (1) applying cross-attention to the state before self-attention, (2) applying cross-attention to the state after self-attention, (3) interleaving cross-attention and self-attention layers, and (4) appending the state as an additional token within the sub-action self-attention block.
(a) CityFlow
(b) 10-D CoNE with 25% pits
Figure 8: Comparison of state conditioning strategies in CityFlow (left) and 10-D CoNE (right). SAINT’s pre-attention FiLM-based conditioning outperforms alternatives more clearly in CityFlow, while all strategies perform similarly in CoNE. Results are averaged over 5 seeds; shaded regions indicate one standard deviation.
As shown in Figure 8, FiLM achieves higher final reward and exhibits more stable learning than the alternatives in CityFlow. In the 10-dimensional CoNE environment, all state conditioning strategies perform comparably, with pre-attention FiLM-based conditioning achieving slightly better final performance. These results suggest that while multiple conditioning mechanisms are viable, FiLM offers an advantage and may contribute to more stable training dynamics.
Appendix D SAINT’s Computation Cost
(a) CityFlow
(b) CoNE
Figure 9: Learning curves comparing SAINT, SAINT with inducing points (SAINT-IP), and the pure factorized baseline in CityFlow and CoNE. Despite higher per-step computation cost, both SAINT and SAINT-IP reach the factored baseline’s final performance much faster and achieve higher final rewards. In CityFlow, SAINT-IP’s policy is worse than SAINT’s, but still outperforms the baseline. Results are averaged over 5 seeds; shaded regions denote one standard deviation.
Figure 9 shows training curves for SAINT, SAINT with inducing points (SAINT-IP), and the factorized baseline in the CityFlow and CoNE environments. While SAINT and SAINT-IP incur higher per-step computational costs due to the Transformer blocks, both methods achieve the factorized baseline’s final performance in less total training time. This reflects their ability to reach performant policies with fewer training episodes. In CityFlow, SAINT-IP exhibits a degradation in asymptotic reward relative to SAINT, but still outperforms the factorized baseline. These results illustrate that the overhead of modeling sub-action dependencies can be offset by more efficient use of training experience.